# Fusion of Multiscale Features Via Centralized Sparse-attention Network for EEG Decoding

### v1

@@ Abstract
Electroencephalography (EEG) signal decoding is a key technology that translates brain activity into executable commands, laying the foundation for direct brain–machine interfacing and intelligent interaction. To address the inherent spatiotemporal heterogeneity of EEG signals, this paper proposes a multi-branch parallel architecture, where each temporal scale is equipped with an independent spatial feature extraction module. To further enhance multi-branch feature fusion, we propose a Fusion of Multiscale Features via Centralized Sparse-attention Network (EEG-CSANet), a centralized sparse-attention network. It employs a main–auxiliary branch architecture, where the main branch models core spatiotemporal patterns via multiscale self-attention, and the auxiliary branch facilitates efficient local interactions through sparse cross-attention. Experimental results show that EEG-CSANet achieves state-of-the-art (SOTA) performance across five public datasets (BCIC-IV-2A, BCIC-IV-2B, HGD, SEED, and SEED-VIG), with accuracies of 88.54%, 91.09%, 99.43%, 96.03%, and 90.56%, respectively. Such performance demonstrates its strong adaptability and robustness across various EEG decoding tasks. Moreover, extensive ablation studies are conducted to enhance the interpretability of EEG-CSANet. In the future, we hope that EEG-CSANet could serve as a promising baseline model in the field of EEG signal decoding. The source code is publicly available at: https://github.com/Xiangrui-Cai/EEG-CSANet.
@@ 1 Introduction
Brain–computer interfaces (BCIs) enable direct communication between the human brain and external devices[1] and have attracted increasing attention due to their broad application potential in human–computer interaction, motor rehabilitation, affective computing, fatigue detection, as well as disease diagnosis and treatment[2][3][4]. Electroencephalography (EEG), which measures voltage fluctuations generated by neural activity, has become the most widely used signal acquisition modality in contemporary BCI systems, owing to its noninvasive nature, high temporal resolution, and strong portability[5]. Currently, the dominant EEG paradigms include motor imagery (MI), event-related P300, and steady-state visual evoked potentials (SSVEP)[6][7]. Owing to their distinct neural response characteristics, these paradigms have demonstrated strong effectiveness and adaptability across a wide range of application scenarios[8][9].
@@
However, efficiently and accurately decoding human intent remains a fundamental challenge in the BCI field, primarily due to the nonstationary nature of EEG signals, substantial inter-subject variability, and low signal noise ratio[10]. To address this issue, a variety of machine learning (ML) and deep learning (DL) based approaches have been proposed to enhance feature extraction and classification performance of brain signals. Traditional ML-based methods typically consist of two sequential stages: feature extraction and classification. During the feature extraction stage, techniques such as Common Spatial Pattern (CSP), Differential Entropy (DE), and Power Spectral Density (PSD) are commonly employed. In the classification stage, classifiers including Linear Discriminant Analysis (LDA) and Support Vector Machine (SVM) have been shown to achieve competitive performance[11][12][13]. Nevertheless, these methods rely heavily on handcrafted feature engineering, which is labor-intensive and highly dependent on domain-specific prior knowledge, thereby limiting their robustness and generalization capability.
@@
By comparison, deep learning based methods employ end-to-end models composed of multiple processing layers, enabling the automatic learning of intrinsic data representations and thereby substantially reducing the reliance on manual intervention and domain-specific preprocessing. With the rapid advancement of deep learning techniques, a wide range of deep learning models have been extensively applied to EEG decoding. Early studies introduced Deep ConvNet, Shallow ConvNet[14], and EEGNet[15], which significantly improved classification accuracy and model robustness, and have since become representative approaches for EEG feature extraction. Subsequently, Long Short-Term Memory (LSTM) networks[16], hybrid CNN–LSTM models[17], and multi-branch CNN architectures[18] were proposed, further advancing EEG decoding performance. In recent years, driven by the success of Transformer architectures in sequence modeling, self-attention mechanisms have been introduced into EEG analysis. Models such as Conformer[19] and ATCNet[20] effectively integrate attention mechanisms with CNN architectures, leading to substantial improvements in feature representation capability and elevating EEG decoding accuracy to a new level. Consequently, models based on the “CNN + Transformer” paradigm have gradually become the dominant approach.
@@
In recent years, multi-scale modeling approaches have gained widespread attention in EEG decoding tasks due to the complex spatiotemporal features of EEG signals[21][22]. Representative works, such as MSTFNet[23], EEGTransNet[24], MCMTNet[25], and TMSA-Net[26], typically employ multi-scale temporal convolutions to extract features at different temporal resolutions, which are subsequently fused through concatenation or addition and then fed into subsequent spatial convolutional layers to capture inter-regional spatial patterns. However, these architectures implicitly rely on a critical assumption: that all temporal scales share the same spatial structure. This assumption contradicts neurophysiological evidence[27][28], which indicates that different temporal scales correspond to distinct frequency components in EEG signals (e.g., $\delta$, $\theta$, $\alpha$, $\beta$, $\gamma$ bands), and that these frequency bands exhibit markedly different patterns of regional activation and functional connectivity in the brain[29][30]. Given the aforementioned heterogeneous spatial correlations across different temporal scales, enforcing a shared spatial modeling mechanism across all scales may overlook frequency-specific spatial topologies, thus limiting the model’s representational capacity and decoding performance.
@@
At the same time, effectively integrating multi-branch features remains a key challenge in feature fusion[31][32]. In this context, the cross-attention mechanism[33] has recently demonstrated strong potential for feature fusion due to its ability to model inter-branch dependencies[34][35]. Cross-attention dynamically captures semantic relationships between different branches through the interaction of queries, keys, and values, and performs information fusion across modalities or hierarchical levels based on attention weights. This mechanism not only strengthens the generalization capability of the learned representations but also facilitates extensive applications in fields such as vision–language processing and multimodal learning[36][37][38].
Motivated by these insights and the studies in[39][40], we propose an innovative fusion architecture that coordinates a primary branch with auxiliary branches. The primary branch employs a multi-scale multi-head self-attention mechanism (MSA) to extract and reinforce core spatiotemporal features, while multiple auxiliary branches utilize a multi-scale sparse multi-head cross-attention mechanism (MSCA) to focus on local details at specific scales. During the fusion stage, MSCA guides the auxiliary branches to interact only with semantically relevant key local regions in the primary branch, effectively avoiding computational redundancy and noise interference caused by global associations. Specifically, multi-scale feature extraction integrates contextual information at different granularities, enriching the overall representation; the cross-attention mechanism ensures precise semantic alignment between primary and auxiliary branches, maintaining consistency between local details and global structure; and the sparsity mechanism further constrains the interaction scope, suppressing irrelevant regions and reducing computational overhead. The synergistic effect of these components enables MSCA to efficiently refine key local features under the semantic guidance of the primary branch, complementing MSA and achieving effective and precise feature aggregation, thereby substantially enhancing the model’s representational capacity and robustness.
@@ 2 Methods
The proposed EEG-CSANet architecture is illustrated in Figure 1. Initially, a same category signal segmentation and reconstruction (S&R)[41] method is employed for data augmentation of EEG signals. The original EEG data are then combined with the augmented data to expand the training sample size. Subsequently, the signals are processed by a multiscale temporal feature extraction module. To avoid losing inter-channel interaction information at different scales, a multi-branch structure is adopted, which separately extracts deep channel features for each time scale. To effectively integrate EEG signals extracted from multiple branches, we have innovatively designed a fusion mechanism: the main branch employs a multiscale multi-head self-attention mechanism, while auxiliary branches introduce a multiscale sparse multi-head cross-attention mechanism, achieving efficient information exchange between branches. The integrated features are then fed into a temporal convolutional network to further extract higher-level temporal dynamic features. To prevent the degradation of temporal information within deep networks, residual structures are incorporated into the convolutional modules. Finally, a fully connected layer is used to complete the classification task, the specific parameters are shown in Table 1. The following sections will provide a detailed description of each component of the EEG-CSANet.
@@ 2.1 Data augmentation
We use the method of S&R to perform data augmentation on EEG signals, which is a commonly used data augmentation method in MI[19][24][25]. Specifically, we let $X_{i}\in\mathbb{R}^{B\times C\times T}$ denote the original EEG signal of the $i$-th class, where $i\in\{1,\dots,L\}$ , $L$ is the total number of classes, and $B$, $C$, and $T$ denote the batch size, number of channels, and number of time samples, respectively. The signal $X_{i}$ is uniformly divided into $S$ segments along the time dimension. New samples $X_{i}^{\prime}$ are then generated by randomly selecting and reorganizing these segments while preserving their original temporal order. This augmentation is applied separately within each training batch, generating an amount of synthetic data equal in size to the original input $X_{i}$. Therefore, the model input for each batch after augmentation is given by:
@@ 2.2 Multi-branch convolution
Previous studies[23][24][25][26] have tended to employ multiscale temporal convolution fusion methods to extract features from EEG signals, as illustrated in Figure 2. This approach concatenates multiscale temporal features to jointly extract channel-wise features. However, we contend that directly concatenating multiscale temporal features for channel-wise feature extraction may not effectively capture these distinct spatial patterns. Therefore, we propose a multi-branch architecture to separately process and fuse features from different temporal scales, as shown in Figure 2. In this architecture, each branch is responsible for handling features at a specific temporal scale or frequency band, thereby enabling more accurate capture and representation of the distinct spatial characteristics present across different frequency bands.
The combined structure of temporal convolution and depthwise separable spatial convolution (DW-Spa-Conv) resembles that of EEGNet[15]. Specifically, the input signal $X_{\mathrm{Input}}$ first passes through a temporal convolutional module with kernel size $(K_{i},1)$, where $i\in\{1,2,3,4\}$ corresponds to the four branches, to extract temporal features for each channel. This is followed by a depthwise separable convolution with kernel size $(C,1)$, where the depth multiplier $D$ controls the number of output feature maps per input channel, enabling spatial feature extraction. Next, the features are downsampled via the first average pooling layer with kernel size $(P_{1},1)$. The pooled features are then fed into a subsequent convolutional module with kernel size $(K_{5},1)$; we regard this module as facilitating further channel-wise interaction, thus enhancing the extraction of high-dimensional spatio-temporal features. Finally, a second average pooling layer with kernel size $(P_{2},1)$ performs additional downsampling. As a result, the output of each of the four convolutional branches is $Z_{i}\in\mathbb{R}^{B\times U_{i}\times T_{0}}$:
@@ 2.3 Feature fusion architecture
Specifically, we designate $Z_{1}$ as the main branch and $Z_{2}$, $Z_{3}$, $Z_{4}$ as auxiliary branches. For the main branch $Z_{1}$, we adopt a relatively large convolutional kernel $K_{1}$ for temporal feature extraction, aiming to capture extensive global spatio-temporal dependencies. In contrast, the auxiliary branches $Z_{2}$, $Z_{3}$, and $Z_{4}$ use smaller kernels $K_{2}$, $K_{3}$, $K_{4}$ to preserve finer local spatio-temporal patterns. Due to the constrained receptive fields of small kernels, they are often inadequate for modeling long-range dependencies and can overlook crucial global structural information. To address this issue, we introduce a multiscale sparse multi-head cross-attention mechanism[39] that enhances global perception while capturing local fine-grained details, enabling collaborative fusion between the main and auxiliary branches.
@@
The attention score matrix $A$ is subsequently computed using scaled dot-product attention[33]:
@@
Before applying the softmax function, we introduce a $\mathrm{Top}\text{-}k$ sparsification operation[39], which discards the smallest 1k1-k proportion of values in each row of $A$. This effectively mitigates the influence of noise from auxiliary branches on feature extraction, as illustrated in Figure 1.D.D. For the softmax operation, these discarded values are replaced with -$\infty$, causing them to be transformed into zero after softmax, similar to a selective masking mechanism. The operation is expressed as:
@@
We apply the $\mathrm{Top}\text{-}k$ operation with two different ratios, controlled by parameters $k_{1}$ and $k_{2}$, and introduce two learnable scalars $\alpha$ and $\beta$ to adaptively balance their contributions, it is a strategy inspired by[40]. Thus, the final attention output is computed as:
@@
This design enables the effective fusion of information from the main branch $X$ and auxiliary branches $Y$, thereby achieving cross-branch feature enhancement. Notably, sparsification is omitted in the main branch, which instead employs a multiscale multi-head self-attention mechanism to retain global temporal patterns. The combination of multiscale pooling and sparsification, however, may result in a significant loss of temporal information. To address this, we adopt a residual connection[42] to retain the original features:

### v2

@@ Abstract
Electroencephalography (EEG) signal decoding is a key technology that translates brain activity into executable commands, laying the foundation for direct brain-machine interfacing and intelligent interaction. To address the inherent spatiotemporal heterogeneity of EEG signals, this paper proposes a multi-branch parallel architecture, where each temporal scale is equipped with an independent spatial feature extraction module. To further enhance multi-branch feature fusion, we propose a Fusion of Multiscale Features via Centralized Sparse-attention Network (EEG-CSANet), a centralized sparse-attention network. It employs a main-auxiliary branch architecture, where the main branch models core spatiotemporal patterns via multiscale self-attention, and the auxiliary branch facilitates efficient local interactions through sparse cross-attention. Experimental results show that EEG-CSANet achieves state-of-the-art (SOTA) performance across five public datasets (BCIC-IV-2A, BCIC-IV-2B, HGD, SEED, and SEED-VIG), with accuracies of 88.54%, 91.09%, 97.15%, 96.03%, and 90.56%, respectively. Such performance demonstrates its strong adaptability and robustness across various EEG decoding tasks. Moreover, extensive ablation studies are conducted to enhance the interpretability of EEG-CSANet. In the future, we hope that EEG-CSANet could serve as a promising baseline model in the field of EEG signal decoding. The source code is publicly available at: https://github.com/Xiangrui-Cai/EEG-CSANet.
@@ 1 Introduction
Brain–computer interfaces (BCIs) enable direct communication between the human brain and external devices[mcfarland2011brain] and have attracted increasing attention due to their broad application potential in human–computer interaction, motor rehabilitation, affective computing, fatigue detection, as well as disease diagnosis and treatment[park2022brain][zhang2023adaptive][edelman2024non]. Electroencephalography (EEG), which measures voltage fluctuations generated by neural activity, has become the most widely used signal acquisition modality in contemporary BCI systems, owing to its noninvasive nature, high temporal resolution, and strong portability[liu2025recent]. Currently, the dominant EEG paradigms include motor imagery (MI), event-related P300, and steady-state visual evoked potentials (SSVEP)[yadav2025decoding][abiri2019comprehensive]. Owing to their distinct neural response characteristics, these paradigms have demonstrated strong effectiveness and adaptability across a wide range of application scenarios[saibene2023eeg][qin2023application].
@@
However, efficiently and accurately decoding human intent remains a fundamental challenge in the BCI field, primarily due to the nonstationary nature of EEG signals, substantial inter-subject variability, and low signal noise ratio[saibene2024deep]. To address this issue, a variety of machine learning (ML) and deep learning (DL) based approaches have been proposed to enhance feature extraction and classification performance of brain signals. Traditional ML-based methods typically consist of two sequential stages: feature extraction and classification. During the feature extraction stage, techniques such as Common Spatial Pattern (CSP), Differential Entropy (DE), and Power Spectral Density (PSD) are commonly employed. In the classification stage, classifiers including Linear Discriminant Analysis (LDA) and Support Vector Machine (SVM) have been shown to achieve competitive performance[ang2012filter][zheng2015investigating][cao2022effective]. Nevertheless, these methods rely heavily on handcrafted feature engineering, which is labor-intensive and highly dependent on domain-specific prior knowledge, thereby limiting their robustness and generalization capability.
@@
By comparison, deep learning based methods employ end-to-end models composed of multiple processing layers, enabling the automatic learning of intrinsic data representations and thereby substantially reducing the reliance on manual intervention and domain-specific preprocessing. With the rapid advancement of deep learning techniques, a wide range of deep learning models have been extensively applied to EEG decoding. Early studies introduced Deep ConvNet, Shallow ConvNet[schirrmeister2017deep], and EEGNet[lawhern2018eegnet], which significantly improved classification accuracy and model robustness, and have since become representative approaches for EEG feature extraction. Subsequently, Long Short-Term Memory (LSTM) networks[zhang2019novel], hybrid CNN–LSTM models[wang20232d], and multi-branch CNN architectures[yang2021two] were proposed, further advancing EEG decoding performance. In recent years, driven by the success of Transformer architectures in sequence modeling, self-attention mechanisms have been introduced into EEG analysis. Models such as Conformer[song2022eeg] and ATCNet[altaheri2022physics] effectively integrate attention mechanisms with CNN architectures, leading to substantial improvements in feature representation capability and elevating EEG decoding accuracy to a new level. Consequently, models based on the “CNN + Transformer” paradigm have gradually become the dominant approach.
@@
In recent years, multi-scale modeling approaches have gained widespread attention in EEG decoding tasks due to the complex spatiotemporal features of EEG signals[zhou2025csbrain][wang2024cbramod]. Representative works, such as MSTFNet[jin2024multiscale], EEGTransNet[ma2024attention], MCMTNet[yang2025mcmtnet], and TMSA-Net[zhao2025tmsa], typically employ multi-scale temporal convolutions to extract features at different temporal resolutions, which are subsequently fused through concatenation or addition and then fed into subsequent spatial convolutional layers to capture inter-regional spatial patterns. However, these architectures implicitly rely on a critical assumption: that all temporal scales share the same spatial structure. This assumption contradicts neurophysiological evidence[li2025frequency][tao2023adfcnn], which indicates that different temporal scales correspond to distinct frequency components in EEG signals (e.g., $\delta$, $\theta$, $\alpha$, $\beta$, $\gamma$ bands), and that these frequency bands exhibit markedly different patterns of regional activation and functional connectivity in the brain[pfurtscheller2006mu][pfurtscheller1997motor]. Given the aforementioned heterogeneous spatial correlations across different temporal scales, enforcing a shared spatial modeling mechanism across all scales may overlook frequency-specific spatial topologies, thus limiting the model’s representational capacity and decoding performance.
@@
At the same time, effectively integrating multi-branch features remains a key challenge in feature fusion[zhao2025deep][han2025multi]. In this context, the cross-attention mechanism[vaswani2017attention] has recently demonstrated strong potential for feature fusion due to its ability to model inter-branch dependencies[cao2025bi][zhao2024feature]. Cross-attention dynamically captures semantic relationships between different branches through the interaction of queries, keys, and values, and performs information fusion across modalities or hierarchical levels based on attention weights. This mechanism not only strengthens the generalization capability of the learned representations but also facilitates extensive applications in fields such as vision–language processing and multimodal learning[li2023blip][jian2024rethinking][shen2024icafusion].
Motivated by these insights and the studies in[chen2023learning][ma2025multiscale], we propose an innovative fusion architecture that coordinates a primary branch with auxiliary branches. The primary branch employs a multi-scale multi-head self-attention mechanism (MSA) to extract and reinforce core spatiotemporal features, while multiple auxiliary branches utilize a multi-scale sparse multi-head cross-attention mechanism (MSCA) to focus on local details at specific scales. During the fusion stage, MSCA guides the auxiliary branches to interact only with semantically relevant key local regions in the primary branch, effectively avoiding computational redundancy and noise interference caused by global associations. Specifically, multi-scale feature extraction integrates contextual information at different granularities, enriching the overall representation; the cross-attention mechanism ensures precise semantic alignment between primary and auxiliary branches, maintaining consistency between local details and global structure; and the sparsity mechanism further constrains the interaction scope, suppressing irrelevant regions and reducing computational overhead. The synergistic effect of these components enables MSCA to efficiently refine key local features under the semantic guidance of the primary branch, complementing MSA and achieving effective and precise feature aggregation, thereby substantially enhancing the model’s representational capacity and robustness.
@@ 2 Methods
The proposed EEG-CSANet architecture is illustrated in Figure 1. Initially, a same category signal segmentation and reconstruction (S&R)[lotte2015signal] method is employed for data augmentation of EEG signals. The original EEG data are then combined with the augmented data to expand the training sample size. Subsequently, the signals are processed by a multiscale temporal feature extraction module. To avoid losing inter-channel interaction information at different scales, a multi-branch structure is adopted, which separately extracts deep channel features for each time scale. To effectively integrate EEG signals extracted from multiple branches, we have innovatively designed a fusion mechanism: the main branch employs a multiscale multi-head self-attention mechanism, while auxiliary branches introduce a multiscale sparse multi-head cross-attention mechanism, achieving efficient information exchange between branches. The integrated features are then fed into a temporal convolutional network to further extract higher-level temporal dynamic features. To prevent the degradation of temporal information within deep networks, residual structures are incorporated into the convolutional modules. Finally, a fully connected layer is used to complete the classification task, the specific parameters are shown in Table 1. The following sections will provide a detailed description of each component of the EEG-CSANet.
@@ 2.1 Data augmentation
We use the method of S&R to perform data augmentation on EEG signals, which is a commonly used data augmentation method in MI[song2022eeg][ma2024attention][yang2025mcmtnet]. Specifically, we let $X_{i}\in\mathbb{R}^{B\times C\times T}$ denote the original EEG signal of the $i$-th class, where $i\in\{1,\dots,L\}$ , $L$ is the total number of classes, and $B$, $C$, and $T$ denote the batch size, number of channels, and number of time samples, respectively. The signal $X_{i}$ is uniformly divided into $S$ segments along the time dimension. New samples $X_{i}^{\prime}$ are then generated by randomly selecting and reorganizing these segments while preserving their original temporal order. This augmentation is applied separately within each training batch, generating an amount of synthetic data equal in size to the original input $X_{i}$. Therefore, the model input for each batch after augmentation is given by:
@@ 2.2 Multi-branch convolution
Previous studies[jin2024multiscale][ma2024attention][yang2025mcmtnet][zhao2025tmsa] have tended to employ multiscale temporal convolution fusion methods to extract features from EEG signals, as illustrated in Figure 2. This approach concatenates multiscale temporal features to jointly extract channel-wise features. However, we contend that directly concatenating multiscale temporal features for channel-wise feature extraction may not effectively capture these distinct spatial patterns. Therefore, we propose a multi-branch architecture to separately process and fuse features from different temporal scales, as shown in Figure 2. In this architecture, each branch is responsible for handling features at a specific temporal scale or frequency band, thereby enabling more accurate capture and representation of the distinct spatial characteristics present across different frequency bands.
The combined structure of temporal convolution and depthwise separable spatial convolution (DW-Spa-Conv) resembles that of EEGNet[lawhern2018eegnet]. Specifically, the input signal $X_{\mathrm{Input}}$ first passes through a temporal convolutional module with kernel size $(K_{i},1)$, where $i\in\{1,2,3,4\}$ corresponds to the four branches, to extract temporal features for each channel. This is followed by a depthwise separable convolution with kernel size $(C,1)$, where the depth multiplier $D$ controls the number of output feature maps per input channel, enabling spatial feature extraction. Next, the features are downsampled via the first average pooling layer with kernel size $(P_{1},1)$. The pooled features are then fed into a subsequent convolutional module with kernel size $(K_{5},1)$; we regard this module as facilitating further channel-wise interaction, thus enhancing the extraction of high-dimensional spatio-temporal features. Finally, a second average pooling layer with kernel size $(P_{2},1)$ performs additional downsampling. As a result, the output of each of the four convolutional branches is $Z_{i}\in\mathbb{R}^{B\times U_{i}\times T_{0}}$:
@@ 2.3 Feature fusion architecture
Specifically, we designate $Z_{1}$ as the main branch and $Z_{2}$, $Z_{3}$, $Z_{4}$ as auxiliary branches. For the main branch $Z_{1}$, we adopt a relatively large convolutional kernel $K_{1}$ for temporal feature extraction, aiming to capture extensive global spatio-temporal dependencies. In contrast, the auxiliary branches $Z_{2}$, $Z_{3}$, and $Z_{4}$ use smaller kernels $K_{2}$, $K_{3}$, $K_{4}$ to preserve finer local spatio-temporal patterns. Due to the constrained receptive fields of small kernels, they are often inadequate for modeling long-range dependencies and can overlook crucial global structural information. To address this issue, we introduce a multiscale sparse multi-head cross-attention mechanism[chen2023learning] that enhances global perception while capturing local fine-grained details, enabling collaborative fusion between the main and auxiliary branches.
@@
The attention score matrix $A$ is subsequently computed using scaled dot-product attention[vaswani2017attention]:
@@
Before applying the softmax function, we introduce a $\mathrm{Top}\text{-}k$ sparsification operation[chen2023learning], which discards the smallest 1k1-k proportion of values in each row of $A$. This effectively mitigates the influence of noise from auxiliary branches on feature extraction, as illustrated in Figure 1.D.D. For the softmax operation, these discarded values are replaced with -$\infty$, causing them to be transformed into zero after softmax, similar to a selective masking mechanism. The operation is expressed as:
@@
We apply the $\mathrm{Top}\text{-}k$ operation with two different ratios, controlled by parameters $k_{1}$ and $k_{2}$, and introduce two learnable scalars $\alpha$ and $\beta$ to adaptively balance their contributions, it is a strategy inspired by[ma2025multiscale]. Thus, the final attention output is computed as:
@@
This design enables the effective fusion of information from the main branch $X$ and auxiliary branches $Y$, thereby achieving cross-branch feature enhancement. Notably, sparsification is omitted in the main branch, which instead employs a multiscale multi-head self-attention mechanism to retain global temporal patterns. The combination of multiscale pooling and sparsification, however, may result in a significant loss of temporal information. To address this, we adopt a residual connection[he2016deep] to retain the original features: